# Dual-Polarized Intelligent Omni-Surfaces for Independent Reflective-Refractive Transmission

### v1

@@ Abstract
Intelligent omni-surface (IOS), which are capable of providing service coverage to mobile users (MUs) in a reflective and a refractive manner, has recently attracted widespread attention. However, the performance of traditionally IOS-aid systems is limited by the intimate coupling between the refraction and reflection behavior of IOS elements. In this letter, to overcome this challenge, we introduce the concept of dual-polarized IOS-assisted communication. More precisely, by employing the polarization domain in the design of IOS, full independent refraction and reflection modes can be delivered. We consider a downlink dual-polarized IOS-aided system, while also accounting for the leakage between different polarizations. To maximize the sum rate, we formulate a joint IOS phase shift and BS beamforming problem and proposed an iterative algorithm to solve the non-convex program. Simulation results validate that dual-polarized IOS significantly enhances the performance than the traditional one.
@@
Dual-polarized intelligent omni surface, phase shift design.
@@ I Introduction
Recently, reconfigurable intelligent surface (RISs)-assisted wireless communication has attracted great attention, since it enables cost-effective and energy-efficient high data rate communication for future sixth-generation (6G) communication systems. An RIS is an ultra-thin surface containing multiple sub-wavelength nearly passive scattering elements, which is capable of adjusting radio propagation conditions by controlling the phase shifts of the signals that impinge on the surface. Intelligent omni-surface (IOS) is an important instance of an RIS. It can provide service coverage to users on both sides of the surface and overcomes the limitation of widely studied RISs, which only serve users on one side of the surface.
@@
An IOS significantly extends the service coverage of the BS when compared to the traditional RIS and shows great potential for 6G communication [1]. In [2], the authors consider a jointly digital and analog beamforming at the base station (BS) and IOS, respectively. In [3], multiple indoor users obtain omni-directionally services from a small base station with the aid of an IOS. In these studies, the power ratio of the reflected and refracted signals is dependent on the structure of each IOS element and cannot be altered once designed [4, 5, 6]. Therefore, when the ratio of users on both sides is unbalanced with the power ratio, the performance of the IOS-assisted communication system is poor. Additionally, the optimal deployment position of the IOS also depends on the fixed power ration [7].
@@
To address this issue and realize the full potential of reconfigurable IOS-aided systems, the concept of dual-polarized IOS is proposed [8]. By utilizing the polarization domain, fully independent refraction and reflection modes can be delivered. The polarized IOS element can treat vertical and horizontal polarizations with different polarization-independent electromagnetic response. Then, the power ratio of the refraction and reflection can be equivalently altered by allocating power to different polarized antennas at the BS.
Figure 1: System model for a dual-polarized IOS-aided MISO downlink system.
@@
Inspired by this, in this letter, we consider a downlink dual-polarized IOS-assisted system, while also accounting for the leakage between different polarizations. The main contributions are summarized as follows:
@@
  • We propose a multi-user dual-polarized IOS-assisted downlink communication system. The physical characteristics of the IOS and the polarization leakage of the considered system model are introduced and discussed
@@
  • Based on this, we formulate a joint BS digital and IOS analog beamforming problem, and aim to maximize the sum rate of the users on both sides of the IOS.
@@
  • We further analyze the impact of the polarization leakage and user ratio on both sides of the IOS on the performance of the dual-polarized IOS-assisted system and compare the system with the traditional IOS-aided one in the simulation.
@@
The rest of this letter is organized as follows. In Section II, we introduce the system model for dual-polarized IOS-assisted communications and formulated the jointly digital and analog beamforming problem. In Section III, the original problem is reformulated as a more tractable form and divided into several sub-problems. Then, an iterative algorithm is proposed for addressing this problem. The theoretical analysis are elaborated in Section IV. Numerical results in Section V validate the performance enhancement of the dual-polarized IOS. Finally, conclusions are drawn in Section VI.
@@ II-A Scenario Description
As shown in Fig. 1, we consider a multi-user polarized multiple-input-multiple-output (MISO) network, where one BS of $N_{t}$ dual-polarized antennas communicates with $2K$ mobile users (MUs). Each user is equipped with an reconfigurable antenna, which allows to select one of the polarizations for receiving signals [8]. The dual-polarized intelligent omni-surface (IOS) with $M$ polarized elements is deployed between the BS and users to reflect or refract the transmit signal. It can choose one polarization for reflection and the other one for refraction. The MUs are divided into two subsets according to their location with respect to the IOS. The subset receiving the reflected signal from the IOS is denoted by ${\mathcal{N}}_{r}$, and the other subset is denoted by ${\mathcal{N}}_{t}$, with ${\mathcal{N}}_{r}\cup{\mathcal{N}}_{t}=2K$.
@@ II-B System Model
The signal transmitted form the BS is expressed as
@@
where ${\bf s}\sim\mathcal{CN}(0,{\bf I})\in\mathbb{C}^{2K\times 1}$ denotes different data streams for the MUs, ${\bf W}=[{\bf W}_{r},{\bf W}_{t}]$ denotes the beamforming matrix, and ${\bf x}_{v}$, ${\bf x}_{h}$ represents the signal transmitted from the vertically and horizontally polarized antennas, respectively.
II-B1 Polarized IOS
The electromagnetic (EM) response $g_{m}$ of the $m^{th}$ IOS element takes different values depending on the signal’s polarization state and are shown as follows.
@@ II-B1 Polarized IOS
The EM response of the $m^{th}$ dual-polarized IOS element to the incident signal, i.e., $g_{m}$, is an angle-dependent value and can be modeled as
@@
We assume that the polarized IOS is equipped with $N$-bits elements. Then, the available phase shift set is denoted by ${\mathcal{S}}_{a}=\{0,...,\frac{2\pi}{2^{N}}(2^{N}-1)\}$ and the phase shift of the $m^{th}$ element $\psi_{m}\in{\mathcal{S}}_{a}$.
It is noted that for dual-polarized IOS, the phase adjustments of the refracted and reflected link can be independently controlled due to the use of the polarization domain in the design and operation of IOS, while for conventional IOS, the refraction and reflection coefficients are intimately coupled due to the electromagnetic properties of the surface [9].
@@ II-B2 Channel Model
In ideal situation, the electromagnetic (EM) waves in orthogonal polarizations exhibit nearly independent propagation characteristics. However, the scatterers in the environment can cause a cross-polarization effect that alters the polarization state of the EM waves and produces polarization leakage. To describe this leakage, we introduce the cross-polarization discrimination (XPD).
@@
where the channel XPD factor is defined as $\beta_{{\rm IU},k}$. The direct path ${\bf h}_{{\rm BU},k}$ between the BS and the $k^{th}$ MU can be similarly modeled.
II-C IOS-based Beamforming
@@
$y_{k}$ $=\left({\bf h}_{{\rm BU},k}+{\bf h}_{{\rm IU},k}{\bf G}{\bf H}_{\rm BI}\right){\bf x}+{n_{k}}$
$={\bf h}_{k}{\bf x}+n_{k},$ (14)
where the IOS coefficient matrix $\bf G$ is denoted by ${\rm diag}\{g^{vv}_{1},...,g^{vv}_{m},g^{hh}_{1},...,g^{hh}_{m}\}$, ${\bf h}_{k}={\bf h}_{{\rm BU},k}+{\bf h}_{{\rm IU},k}{\bf G}{\bf H}_{\rm BI}$ represents the equivalent channel between the BS and the $k^{th}$ user, and $n_{k}$ is the additive white Gaussian noise (AWGN) at the $k^{th}$ user whose mean is zero and variance is $\sigma^{2}$.
Assume that the channel state information (CSI) is known at the transmitter. Based on the received signal in (II-C), the signal-to-interference-plus-noise ratio (SINR) of the $k^{th}$ MU in the reflected side can be formulated as
@@
where ${\bf W}^{(k)}$ denotes the $k^{th}$ column of the beamformer $\bf W$, $\sum_{k_{r}^{\prime}\neq k_{r}}|{\bf h}_{r,k_{r}}{\bf W}_{r}^{(k_{r}^{\prime})}|^{2}$ denotes the interference of MUs with same polarization, and $\sum_{i\in{\mathcal{N}}_{t}}|{\bf h}_{r,k_{r}}{\bf W}_{t}^{(i)}|^{2}$ represents the polarization leakage of other MUs with different polarization. The SINRs of the users equipped with a horizontally polarized antenna can be similarly written.
II-D Problem formulation
@@
where $P_{\rm BS}$ denotes the maximum transmit power at the BS.
@@ III-A Reformulate of the original problem
The minimum mean-square error (MMSE) algorithm is employed for addressing the non-convex (16a). By introducing the auxiliary variables ${\bf f}=[f_{r,1},...,f_{r,N_{r}},f_{t,1}...,f_{t,N_{t}}]^{\rm T}$, and ${\bf u}=[u_{r,1},...,u_{r,N_{r}},u_{t,1}...,u_{t,N_{t}}]^{\rm T}$, Problem (16) can be reformulated as
@@
$\operatorname{e}_{\mathrm{r},k_{\mathrm{r}}}=$ $\mathbb{E}\left\{\left({u}_{\mathrm{r},k_{{r}}}^{*}y_{\mathrm{r},k_{\mathrm{r}}}-s_{\mathrm{r},k_{\mathrm{r}}}\right)\left(u_{\mathrm{r},k_{\mathrm{r}}}^{*}y_{\mathrm{r},k_{\mathrm{r}}}-s_{\mathrm{r},k_{\mathrm{r}}}\right)^{*}\right\}$
$=$ $\sum_{i\in\mathcal{K}_{\mathrm{r}}}\left|u_{\mathrm{r},k_{\mathrm{r}}}^{*}{\bf h}_{r,k_{r}}\mathbf{W}^{(i)}_{\mathrm{r}}\right|^{2}+\sum_{j\in\mathcal{K}_{\mathrm{t}}}\left|u_{\mathrm{r},k_{\mathrm{r}}}^{*}{\bf h}_{r,k_{r}}\mathbf{W}^{(j)}_{\mathrm{t}}\right|^{2}$
$-2\rm{Re}\left\{u_{\mathrm{r},k_{\mathrm{r}}}^{*}{\bf h}_{r,k_{r}}\mathbf{W}^{(k_{r})}_{\mathrm{r}}\right\}+\left|u_{\mathrm{r},k_{\mathrm{r}}}\right|^{2}\sigma^{2}+1.$ (18)
By taking the first derivative of (17a) and (18) equal to zero, we obtain the optimal solutions of $f_{r,k_{r}}$ and $u_{r,k_{r}}$ as
@@
III-B Digital beamforming optimzation at the BS
@@
where (·) denotes the matrix pseudo-inverse. The value of $\lambda$ should satisfy the complementary slackness condition $\lambda(\sum_{i\in{\mathcal{N}}_{r}}{\bf W}_{r}^{(i)\rm H}{\bf W}_{r}^{(i)}+\sum_{j\in{\mathcal{N}}_{t}}{\bf W}_{t}^{(j)\rm H}{\bf W}_{t}^{(j)}-P_{\rm BS})=0$. The value of $\lambda$ should be chosen for ensuring the following equation:
@@
III-C Analog beamforming optimzation at the IOS
Given the optimal $\bf f$, $\bf u$, ${\bf W}_{r}$, and ${\bf W}_{t}$, the discrete analog beamforming subproblem can be written as
@@ III-C Analog beamforming optimzation at the IOS
Problem (26) is tackled in two steps: (i) we first consider a relaxed problem with continuous phase shifts and (ii) then a branch-and-bound based algorithm is proposed to account for the discrete phase shifts.
The objective function in (26) is a convex function of the IOS coefficient matrix $\bf G$. By releasing the non-convex (16c), it can be solved by standard tools such as CVX.
However, $\psi^{{\mathcal{X}},opt}_{m}$ lies in the range of $l_{m}\Delta\psi$ and $(l_{m}+1)\Delta\psi$ , where $l_{m}$ denotes an integer satisfying $0\leq l_{m}\Delta\psi\leq\frac{2\pi}{2^{N}}(2^{N}-1)$, and $\Delta\psi=\frac{2\pi}{2^{N}}$ represents the discrete phase shift step, which can be addressed by branch-and-bound (BnB) algorithm [3].
@@ III-D Overall Algorithm
[…]
5:   Calculate the $\lambda$ by adopting bisection search method in $[0,\lambda^{ub})$ based on (25);
6:   Update ${\bf W}_{r}$ and ${\bf W}_{t}$ by (23)-(24);
7:   Update $\bf G$ by solving Problem (26) relaxing (16c);
8:   Calculate the optimal $\bf G$ with discrete phase shift by adopting BnB algorithm;
9:  Until the objective function (17a) converges.
IV Performance Analysis of the IOS-assisted communication system
In this section, we first discuss the impact of the XPD on the system performance, and then analyze the downlink sum rate as a function of the ratio of the MUs.

### v2

# Exploiting Polarization Domain of the IOS for Enhanced Full-Dimensional Transmission

@@ Abstract
Intelligent omni-surface (IOS), capable of providing service coverage to mobile users (MUs) in a reflective and refractive manner, has recently attracted widespread attention. However, the performance of power-domain IOS-assisted systems is limited by the intimate coupling between the refraction and reflection behavior of IOS elements. In this paper, we introduce the concept of dual-polarized IOS-assisted communication to overcome this challenge. By employing the polarization domain in the design of IOS, full independent refraction and reflection modes can be delivered. We consider a downlink dual-polarized IOS-aided system while also accounting for the leakage between different polarizations. To maximize the sum rate, we formulated a joint base station (BS) digital and IOS analog beamforming problem and proposed an iterative algorithm to solve the non-convex program. Simulation results validate that dual-polarized IOS significantly enhances the performance than that of the power-domain one.
@@
Dual-polarized intelligent omni-surface, phase shift design.
@@ I Introduction
Recently, Intelligent omni-surface (IOS)-assisted wireless communication has attracted great attentions since it enables cost-effective and energy-efficient high data rate communication for future sixth-generation (6G) communication systems [1]. It can provide service coverage to users on both sides of the surface and overcomes the limitation of widely studied reconfigurable intelligent surfaces (RISs), which only serve users on one side of the surface [2, 3].
@@
There has been extensive research on IOSs, demonstrating its potential to significantly enhance the service coverage of a base station (BS) compared to traditional RISs. In [4], multiple indoor users obtain omni-directionally services from a small base station with the aid of an IOS. In [5], the authors consider joint digital and analog beamforming at the BS and IOS, respectively. In these studies, the power ratio of the reflected and refracted signals is dependent on the structure of the IOS elements and cannot be altered once designed [6, 7]. Therefore, when the ratio of users on both sides is unbalanced with the power ratio, the performance of the power-domain IOS-assisted communication system deteriorates.
@@
To address this issue and realize the full potential of IOS-aided systems, the concept of dual-polarized IOS is proposed [8]. By utilizing the polarization domain, fully independent refraction and reflection modes can be delivered. The polarized IOS element can treat vertical and horizontal polarizations with polarization-independent electromagnetic (EM) responses. Then, the power ratio of the refraction and reflection can be equivalently altered by allocating power to different polarized antennas at the BS.
@@
Inspired by this, in this paper, we consider a downlink dual-polarized IOS-assisted system, while also accounting for the leakage between different polarizations. This polarization leakage has a significant impact on the system’s performance, and a thorough analysis of the effects it causes is necessary. The main contributions are summarized as follows:
@@
  • We propose a multi-user dual-polarized IOS-assisted downlink communication system. The physical characteristics of the IOS and its polarization leakage are introduced and discussed.
@@
  • Based on this, we formulate a joint BS digital and IOS analog beamforming problem and aim to maximize the sum rate of the users on both reflective and refractive sides of the dual-polarized IOS.
@@
  • We further analyze the impact of the polarization leakage and user ratio on both sides of the IOS on the performance of the dual-polarized IOS-assisted system and compare the system with the power-domain IOS-aided one in the simulation.
@@
The rest of this paper is organized as follows. In Section II, we introduce the system model for dual-polarized IOS-assisted communications and formulate the jointly digital and analog beamforming problem. In Section III, the original problem is reformulated as a more tractable form and divided into several sub-problems. Then, an iterative algorithm is proposed for addressing this problem. The theoretical analysis is elaborated in Section IV. Numerical results in Section V validate the performance enhancement of the dual-polarized IOS. Finally,conclusions are drawn in Section VI.
Figure 1: System model for a dual-polarized IOS-aided downlink system.
@@ II-A Scenario Description
As shown in Fig. 1, we consider a multi-user network where a BS equipped with $N_{t}$ dual-polarized antennas communicates with $2K$ mobile users (MUs). Each MU is equipped with a reconfigurable antenna that can select one of the two polarizations for signal reception [8]. A dual-polarized IOS comprising $M$ polarized elements is deployed between the BS and the MUs to reflect or refract transmitted signals, with the capability to select one polarization for reflection and the other for refraction.
The set of all MUs is partitioned into two subsets based on their relative positions to the IOS. The subset ${\mathcal{N}}_{r}$ contains $K_{r}$ MUs that receive signals reflected by the IOS, while the subset ${\mathcal{N}}_{t}$ contains $K_{t}$ MUs that receive signals refracted by the IOS. The total number of users satisfies ${\mathcal{N}}_{r}\cup{\mathcal{N}}_{t}=2K$.
@@ II-B System Model
The signals transmitted from the BS are expressed as
@@
where ${\bf s}\sim\mathcal{CN}(0,{\bf I})\in\mathbb{C}^{2K\times 1}$ denotes different data streams for the MUs, ${\bf W}=[{\bf W}_{r},{\bf W}_{t}]$ denotes the digital beamforming matrix, and ${\bf x}_{v}$, ${\bf x}_{h}$ represents the signal transmitted from the vertically and horizontally polarized antennas, respectively.
II-B1 Dual-polarized IOS
The EM response $g_{m}$ of the $m^{th}$ IOS element takes different values depending on the signal’s polarization state and is given by
@@ II-B1 Dual-polarized IOS
Furthermore, $g_{m}$ is an angle-dependent value and can be modeled as
@@
We assume that the dual-polarized IOS has $N$-bits quantized phase shifts. Then, the available phase shift set is denoted by ${\mathcal{S}}_{a}=\{0,...,\frac{2\pi}{2^{N}}(2^{N}-1)\}$ and the phase shift of the $m^{th}$ element $\psi_{m}\in{\mathcal{S}}_{a}$.
It is noted that for the dual-polarized IOS, the phase adjustments of the refracted and reflected link can be independently controlled due to the use of the polarization domain of the dual-polarized IOS, while for power-domain IOS, the refraction and reflection coefficients are intimately coupled due to the EM properties of the surface [9].
@@ II-B2 Channel Model
In an ideal situation, the EM waves in orthogonal polarizations exhibit nearly independent propagation characteristics. However, the scatterers in the environment can cause a cross-polarization effect that alters the polarization state of the EM waves and then produces polarization leakage, which can be measured by cross-polarization discrimination (XPD).
@@
where the XPD factor is defined as $\beta_{{\rm IU},k}$. The direct path ${\bf h}_{{\rm BU},k}$ between the BS and the $k^{th}$ MU is modeled likewise.
II-C Problem formulation
@@
$y_{k}={\bf h}_{k}{\bf x}+n_{k},$ (14)
where $n_{k}$ is the additive white Gaussian noise (AWGN) at the $k^{th}$ user whose mean is zero and variance is $\sigma^{2}$. The overall channel between the BS and the $k^{th}$ user is written as
${\bf h}_{k}={\bf h}_{{\rm BU},k}+{\bf h}_{{\rm IU},k}{\bf G}{\bf H}_{\rm BI},$ (15)
where the IOS coefficient matrix $\bf G$ is denoted by ${\rm diag}\{g^{vv}_{1},...,g^{vv}_{m},g^{hh}_{1},...,g^{hh}_{m}\}$.
Assume that the channel state information (CSI) is known at the BS. Based on the received signal in (14), the signal-to-interference-plus-noise ratio (SINR) of the $k_{r}\in{\mathcal{N}}_{r}$ MU in the reflective side can be formulated as
@@
where ${\bf W}^{(k)}$ denotes the $k^{th}$ column of the beamformer $\bf W$, $\sum_{k_{r}^{\prime}\neq k_{r}}|{\bf h}_{r,k_{r}}{\bf W}_{r}^{(k_{r}^{\prime})}|^{2}$ denotes the interference of MUs with same polarization, and $\sum_{i\in{\mathcal{N}}_{t}}|{\bf h}_{r,k_{r}}{\bf W}_{t}^{(i)}|^{2}$ represents the polarization leakage of other MUs with different polarization. The SINRs of the MUs on the refractive side of the IOS can be similarly written.
@@
where $P_{\rm BS}$ denotes the maximum transmit power allowed by the BS.
@@ III-A Reformulate of the original problem
The minimum mean-square error (MMSE) algorithm is employed for addressing the non-convex (17a). By introducing the auxiliary variables ${\bf f}=[f_{r,1},...,f_{r,K_{r}},f_{t,1}...,f_{t,K_{t}}]^{\rm T}$, and ${\bf u}=[u_{r,1},...,u_{r,K_{r}},u_{t,1}...,u_{t,K_{t}}]^{\rm T}$, Problem (17) can be reformulated as
@@
$\operatorname{e}_{{r},k_{{r}}}=$ $\mathbb{E}\left\{\left({u}_{{r},k_{{r}}}^{*}y_{{r},k_{{r}}}-s_{{r},k_{{r}}}\right)\left(u_{{r},k_{{r}}}^{*}y_{{r},k_{{r}}}-s_{{r},k_{{r}}}\right)^{*}\right\}$
$=$ $\sum_{i\in\mathcal{N}_{{r}}}\left|u_{{r},k_{{r}}}^{*}{\bf h}_{r,k_{r}}\mathbf{W}^{(i)}_{{r}}\right|^{2}+\sum_{j\in\mathcal{N}_{{t}}}\left|u_{{r},k_{{r}}}^{*}{\bf h}_{r,k_{r}}\mathbf{W}^{(j)}_{{t}}\right|^{2}$
$-2{\rm{Re}}\left\{u_{{r},k_{{r}}}^{*}{\bf h}_{r,k_{r}}\mathbf{W}^{(k_{r})}_{{r}}\right\}+\left|u_{{r},k_{{r}}}\right|^{2}\sigma^{2}+1.$ (19)
By taking the first derivative of (18a) equal to zero, we obtain the optimal solutions of $f_{r,k_{r}}$ and $u_{r,k_{r}}$ as
@@
III-B Digital beamforming optimization at the BS
@@
where (·) denotes the matrix pseudo-inverse. The value of $\lambda$ should satisfy the complementary slackness condition $\lambda(\sum_{i\in{\mathcal{N}}_{r}}{\bf W}_{r}^{(i)\rm H}{\bf W}_{r}^{(i)}+\sum_{j\in{\mathcal{N}}_{t}}{\bf W}_{t}^{(j)\rm H}{\bf W}_{t}^{(j)}-P_{\rm BS})=0$. The value of $\lambda$ should be chosen to ensure the following equation:
@@
III-C Analog beamforming optimization at the IOS
Given $\bf f$, $\bf u$, ${\bf W}_{r}$, and ${\bf W}_{t}$, the discrete IOS analog beamforming subproblem can be written as
@@ III-C Analog beamforming optimization at the IOS
The objective function in (27) is a convex function of the IOS coefficient matrix $\bf G$. By releasing the non-convex (17c), it can be solved by standard tools such as CVX. However, $\psi^{{\mathcal{X}},opt}_{m}$ lies in the range of $l_{m}\Delta\psi$ and $(l_{m}+1)\Delta\psi$, where $l_{m}$ denotes an integer satisfying $0\leq l_{m}\Delta\psi\leq\frac{2\pi}{2^{N}}(2^{N}-1)$, and $\Delta\psi=\frac{2\pi}{2^{N}}$ represents the discrete phase shift step, which can be addressed by the BnB based algorithm [4].
@@ III-D Overall Algorithm
4:   Calculate $\bf G$ by solving Problem (27) relaxing (17c);
5:   Calculate the optimal $\bf G$ with discrete phase shifts by adopting BnB algorithm;
6:   Calculate ${\bf W}$ with the Lagrangian multiplier method;
7:  Until the objective function (18a) converges.
We denote the sum-rate at the $i^{th}$ iteration by $R({\bf W}^{i},{\bf G}^{i})$. At the $(i+1)^{th}$ iteration, the solution of the optimal IOS analog beamforming given by ${\bf W}^{i}$ yields a sum-rate $R({\bf W}^{i},{\bf G}^{i+1})\geq R({\bf W}^{i},{\bf G}^{i})$. Similarly, the solution of the optimal BS digital beamforming given by ${\bf G}^{i+1}$ yields $R({\bf W}^{i+1},{\bf G}^{i+1})\geq R({\bf W}^{i},{\bf G}^{i+1})$. Therefore, we obtain the following inequalities $R({\bf W}^{i+1},{\bf G}^{i+1})\geq R({\bf W}^{i},{\bf G}^{i})$. In addition, due to the total transmit power constraint, the sum rate has an upper bound. Hence, Algorithm 1 is guaranteed to converge.
The main complexity of Step 6 lies in calculating $\mathbf{W}^{(k)}$ in (24)-(25), especially in the matrix inversion of ${\bf M}\in{\mathbb{C}}^{2N_{t}\times 2N_{t}}$. Therefore, the complexity is in the order of $\mathcal{O}(KN_{t}^{3})$. The complexity of analog beamforming problem is the sum of the continuous and discrete phase shift subproblems. The continuous one lies in the interior point method and can be referred to in [3]. Hence, it is with the complexity of ${\mathcal{O}}(M^{4.5})$, and the order of the BnB algorithm is, in the worst case, $\mathcal{O}({M^{4.5}2^{M}})$ [4]. Thus, the total complexity is $\mathcal{O}({KN_{t}^{3}+M^{4.5}2^{M}})$.
IV Performance Analysis of the Dual-polarized IOS-assisted communication system